OpenStudy (anonymous):

find the distance between (0 , -3) and (8, 9)

OpenStudy (unklerhaukus):

\[√(8^2+12^2)\]

OpenStudy (callisto):

distance \[=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\] \[=\sqrt{(0-8)^2+(-3-9)^2}\] \[=\sqrt{(-8)^2+(-12)^2}\]\[=\sqrt{64+144}\] Can you do it now?

OpenStudy (anonymous):

yes i figured it out ! thank you!

OpenStudy (callisto):

welcome :)

OpenStudy (anonymous):

\[(0,\sqrt{5}) and 3\sqrt{5} , 0\]

OpenStudy (anonymous):

need help with that one

OpenStudy (anonymous):

ya callisto is right ;-)

OpenStudy (callisto):

put (x1,y1) = (0, sqrt5) and (x2,y2)=(3sqrt5,0) into the distance formula, what do you get?

OpenStudy (anonymous):

haha thanks